# A Modified Cosmic Brane Proposal for Holographic Renyi Entropy

### v2

@@ 5 An Illustrative Example: Gaussian Distribution
Figure 5: The Renyi entropy (left) and refined Renyi entropy (right) computed by the modified cosmic brane proposal at a fixed generic value of $n<1$ is shown as a function of $A_{1,0}$ holding all other parameters fixed, along with the answer predicted by the cosmic brane proposal. The three phases (1, 2 and D) arising as we move across the entanglement phase transition are labelled.
Note that the existence of these corrections to the cosmic brane proposal does not rely on being close to an entanglement phase transition, although the nearer we are to a phase transition, the corrections to the cosmic brane proposal arise closer to $n=1$. For illustration, in Fig. 5, we depict the different phases that arise for a fixed generic value of $n<1$ as we increase $A_{1,0}$ across an entanglement phase transition, holding everything else fixed. The Renyi entropy at $n<1$ is given by
@@
and the refined Renyi entropy follows the same pattern of phases. The results are plotted in Fig. 5. Since $\sigma=O(\sqrt{G})$ for gravitational states prepared by a smooth path integral, it is clear from this example that there is an $O(1)$ window of areas around the entanglement phase transition, where there are $O\left(\frac{1}{G}\right)$ corrections to the cosmic brane proposal.
While we only discussed the simplest example where we can observe leading order corrections to the cosmic brane proposal, the results for an arbitrary Gaussian are discussed in detail in Appendix A.
@@ 6.1 PSSY Model
Figure 7: The Renyi entropy (left) and refined Renyi entropy (right) computed by the modified cosmic brane proposal at a generic value of $n<1$ is shown as a function of $\log k$ holding all other parameters fixed, along with the answer predicted by the cosmic brane proposal. The three phases (1, D and 2) arising as we move across the entanglement phase transition are labelled.
@@
As expected, this is consistent with the cosmic brane proposal since Phase D never appears.
@@
Both of these results are plotted in Fig. 7 to contrast with the cosmic brane proposal.
@@ 7 Discussion
To summarize, we have offered a modified cosmic brane proposal to compute Renyi entropies in holographic systems that should be interpreted as an update to the Lewkowycz-Maldacena proposal Lewkowycz:2013nqa . Our modified proposal reproduces all previously known results for the holographic Renyi entropy: it always agrees with the cosmic brane proposal for $n\geq 1$ and explains leading order corrections to the $n<1$ Renyi entropy found in various situations. We now comment on various aspects of our work and possible future directions.

### v3

@@ 5 An Illustrative Example: Gaussian Distribution
Figure 5: The Renyi entropy (left) and refined Renyi entropy (right) computed by the modified cosmic brane proposal at a fixed generic value of $n<1$ is shown as a function of $A_{1,0}$ holding all other parameters fixed, along with the answer predicted by the original cosmic brane proposal. The three phases (1, 2 and D) arising as we move across the entanglement phase transition are labelled.
Note that the existence of these corrections to the original cosmic brane proposal does not rely on being close to an entanglement phase transition, although the nearer we are to a phase transition, the corrections to the original cosmic brane proposal arise closer to $n=1$. For illustration, in Fig. 5, we depict the different phases that arise for a fixed generic value of $n<1$ as we increase $A_{1,0}$ across an entanglement phase transition, holding everything else fixed. The Renyi entropy at $n<1$ is given by
@@
and the refined Renyi entropy follows the same pattern of phases. The results are plotted in Fig. 5. Since $\sigma=O(\sqrt{G})$ for gravitational states prepared by a smooth path integral, it is clear from this example that there is an $O(1)$ window of areas around the entanglement phase transition, where there are $O\left(\frac{1}{G}\right)$ corrections to the original cosmic brane proposal.
While we only discussed the simplest example where we can observe leading order corrections to the original cosmic brane proposal, the results for an arbitrary Gaussian are discussed in detail in Appendix A.
@@ 6.1 PSSY Model
Figure 7: The Renyi entropy (left) and refined Renyi entropy (right) computed by the modified cosmic brane proposal at a generic value of $n<1$ is shown as a function of $\log k$ holding all other parameters fixed, along with the answer predicted by the original cosmic brane proposal. The three phases (1, D and 2) arising as we move across the entanglement phase transition are labelled.
@@
As expected, this is consistent with the original cosmic brane proposal since Phase D never appears.
@@
Both of these results are plotted in Fig. 7 to contrast with the original cosmic brane proposal.
@@ 7 Discussion
To summarize, we have offered a modified cosmic brane proposal to compute Renyi entropies in holographic systems that should be interpreted as an update to the Lewkowycz-Maldacena proposal Lewkowycz:2013nqa . Our modified proposal reproduces all previously known results for the holographic Renyi entropy: it always agrees with the original cosmic brane proposal for $n\geq 1$ and explains leading order corrections to the $n<1$ Renyi entropy found in various situations. We now comment on various aspects of our work and possible future directions.